# Shortcuts Everywhere and Nowhere: Exploring Multi-Trigger Backdoor Attacks

### v2

@@ 4.4 Re-evaluating Existing Defenses
Backdoor Model Detection Figure 7(c) presents the detection AUROC of 4 advanced backdoor model detection methods, including NC, MMBD, UMD, and RNP-U. As can be observed, the four detection methods perform well on All2One multi-trigger attacks, yet perform poorly on All2All and All2Random attacks. Specifically, the detection AUROC deteriorates by nearly 50% and 40% on All2All and All2Random attacks respectively, when compared to that on All2One attacks. Such a huge degradation challenges the reliability of the four detection models under more complex scenarios with unfixed adversarial targets. Out of the four methods, RNP-U consistently outperforms the other three methods, which validates the usefulness of unlearning the benign features before detecting the backdoor features. Meanwhile, although these detection methods can detect the backdoor models to some extent, they all fail to identify the backdoor labels of All2All and All2Random attacks.
@@
Backdoor Removal Table 2 reports the defense performance of four backdoor removal methods, including FT, FP, NAD, and ANP. We measure the defense performance by the remaining ASR, the lower the better. None of the defense methods can fully remove the multi-trigger from the model. The average ASRs are still well above 50%, except for NAD and ANP against All2One attacks. Compared to NAD and ANP, fine-tuning and fine-pruning methods FT and FP are essentially ineffective against multi-trigger attacks, with an ASR $>60\%$. This is because the presence of multiple triggers prevents effective forgetting of the backdoor correlation during standard fine-tuning. Additionally, the coexistence of multiple triggers leads to inconsistent parameter activations, rendering the activation-based pruning method FP ineffective. The current state-of-the-art defense method ANP performs poorly on All2All and All2Random attacks, limiting its practicability against more realistic attacks that have different target labels. Clean accuracy (CA) is another perspective looking into the effectiveness of the defense method. The four backdoor removal methods cause a significant reduction in the model’s clean accuracy when facing multi-trigger attacks. Specifically, against All2One attacks, FP, NAD, and ANP reduce the model’s CA by 10.98%, 8.48%, and 8.68% respectively, significantly impairing the functionality of the model. Based on these results, we conclude that existing backdoor detection and removal methods struggle to address the threat of multi-trigger attacks, and developing more advanced countermeasures is imperative. More results of the removal performance for ViT models are deferred into appendix.
5 Conclusion
@@
Appendix A Experimental Details
All experiments were run on NVIDIA Tesla A100 GPUs with PyTorch implementations. We trained the backdoored models of the CNN architecture from scratch for 60 epochs, using SGD optimizer, batch size 128, weight decay $5\times 10^{-4}$, and an initial learning rate of 0.1 which was decayed to 0.01 at the 40th epoch. For the Transformer architecture, we fine-tuned the model with publicly available pre-trained weights [^1] https://github.com/huggingface/pytorch-image-models for 5 to 10 epochs, while injecting the backdoor with a fine-tuning (learning) rate of 0.001 (other parameters were kept unchanged).
A.1 Datasets and Models
The datasets and DNN models used in our experiments are summarized in Table 3.
Table 3: Datasets and models used in our experiments.
A.2 Attack Configurations
On the CIFAR-10 dataset, we considered 10 prevalent poisoning-based backdoor attack methods, including static triggers like OnePixel (Tran et al., 2018), BadNets (Gu et al., 2017), Trojan attack (Liu et al., 2018b), global triggers like Blend (Chen et al., 2017), Sinusoidal Spectrum (SIG) (Barni et al., 2019), Adversarial Noise (Adv) (Turner et al., 2019), Smooth (Zeng et al., 2021), Nashivell filter (Nash) (Liu et al., 2019), and sample-wise triggers like Dynamic (Nguyen & Tran, 2020), WaNet (Nguyen & Tran, 2021). Figure 8 illustrates 10 types of backdoor triggers and the basic configurations used by these attacks are reported in Table 4. To ensure fair and consistent comparison with previous works, we employed the dirty-label poisoning setting, which involves adding backdoor triggers and modifying the ground truth labels. Due to a fail of reproduction, on the ImageNet dataset, we only considered 5 attacks: BadNets (Gu et al., 2017), Trojan (Liu et al., 2018b), Blend (Chen et al., 2017), and Sinusoidal Spectrum (SIG) (Barni et al., 2019), and Nash (Liu et al., 2019). All backdoored models were trained with standard data augmentation techniques including random crop and horizontal flipping.
Figure 8: The 10 types of backdoor triggers on CIFAR-10.
To ensure the independent effectiveness of the 10 trigger patterns, we validated the attack success rate (ASR) and accuracy (CA) of each trigger on the CIFAR-10 dataset. The experimental results are shown in Table 5, demonstrating that under the single-trigger setting, all attacks reached full (100%) ASR.
Table 4: Configurations of the 10 basic backdoor attacks.
Attacks Trigger Type Trigger Size Trigger Pattern OnePixel Static Pixel Pixel modification BadNets Static Patch checkerboard Trojan Static Patch Reversed watermark Blend Global Full image Gaussian noise SIG Global Full image Sinusoidal signal Adv Sample-wise Full image Noise + checkerboard Smooth Global Full image Style transfer Nash Global Full image Instagram filter Dynamic Sample-wise Full image Mask generator WaNet Sample-wise Full image Deformation
Table 5: The ASR (%) of single-trigger attacks on ResNet-18 and CIFAR-10 dataset. The best results are boldfaced.
Poisoning Rate Single-trigger Attacks CA ASR Attack 1% (500) 0.1% (50) 0.05% (25) 0.02% (10) 1% (500) 0.1% (50) 0.05% (25) 0.02% (10) OnePixel 92.69 92.09 94.79 91.69 94.62 44.23 5.37 1.75 BadNets 93.7 92.79 93 91.6 100 100 11.3 1.1 Trojan 92.99 94.41 91.99 92.99 98.13 16.35 1.53 1.31 Blend 93.7 91.6 94.49 90.99 100 99.88 96.33 84.97 SIG 93 93.29 93.39 91.4 100 96.71 91.09 62.81 Adv 94.19 93.99 91.89 93.49 100 99.1 30.28 0.99 Smooth 91.69 92.6 92.99 92.99 94.9 34.15 16.37 5.84 Nash 93.69 93.39 93.79 91.69 96.18 37.1 12.22 4.15 Dynamic 92.99 94.59 95.19 92.19 99.88 95.94 94.19 19.46 WaNet 93.51 92.89 93.67 93.06 99.23 97.77 59.61 6.98 Average 93.22 93.16 93.52 92.21 98.29 72.12 41.83 18.94
Table 6: The ASR (%) of parallel multi-trigger attacks on the ImageNet-20 subset. The best results are boldfaced.
A.3 Defense Configurations
We considered 8 advanced backdoor defense methods, including 4 backdoored model detection techniques: Neural Cleanse (NC) (Wang et al., 2019), UMD (Xiang et al., 2023), MMDB (Wang et al., 2023), and Unlearning (Li et al., 2023), as well as 4 mainstream backdoor removal techniques, including Fine-tuning (FT), Finepruning (FP) (Liu et al., 2018a), Neural Attention Distillation (NAD) (Li et al., 2021c), and Adversarial Neural Pruning (ANP) (Wu & Wang, 2021). In the case of All defenses had limited access to only 500 clean defense samples.
For the 4 backdoored model detection methods, we used their open-source code and recommended settings. For NC (Wang et al., 2019), we set the learning rate to 0.005 for reverse-engineering the trigger. For MMBD (Wang et al., 2023), we estimated the maximum margin statistic for each category, calculated the $p$-value, and set the threshold to 0.05 to determine the backdoor model. For UMD (Xiang et al., 2023), we set the learning rate to $1e^{-5}$ to optimize the perturbed images and set $\beta=0.05$ to compute the threshold. For RNP-U (Li et al., 2023), we unlearned the model with a learning rate of 0.005 and judged the backdoor model with prediction labels.
For the 4 backdoor removal methods, we used their open-source code and carefully adjusted hyperparameters to achieve the best defense performance. For FT, we fine-tuned the model for 10 epochs with a learning rate of 0.05, 0.01, and 0.001. We re-implemented FP (Liu et al., 2018a) using PyTorch and pruned the last convolutional layer (i.e., Layer4.conv2) of the model. For NAD (Li et al., 2021c), we cautiously selected the best hyperparameter $\beta$ from $[0,5000]$. For ANP (Wu & Wang, 2021), we used their open-source code with the recommended settings, with the perturbation budget set to $\epsilon=0.4$ and the trade-off coefficient to $\alpha=0.2$. We determined the optimal pruning threshold based on grid search. All the above defense techniques used the same data augmentation techniques as model training, i.e., random crop and horizontal flipping.
Table 7: The defense performance (remaining ASR, %) of fine-tuning (FT) against the parallel MTBAs on ViT-base model at poisoning rate 10% (each trigger poisons 1%). The experiments were done on CIFAR-10 using only 1% (500) clean samples as the defense data.
No Defense FT (lr=0.05) FT (lr=0.01) FT (lr=0.001) Backdoor Removal All2One All2All All2Random All2One All2All All2Random All2One All2All All2Random All2One All2All All2Random Clean Acc. 98.68 98.44 98.50 98.60 90.44 98.00 96.46 98.20 97.50 98.72 98.34 98.20 OnePixel 93.86 84.40 84.71 93.91 82.40 78.65 85.96 60.60 75.29 90.85 76.20 91.95 BadNets 100.00 97.20 52.87 100.00 93.40 59.78 100.00 89.20 50.57 99.77 93.80 52.69 Trojan 100.00 95.20 98.90 100.00 94.60 98.81 100.00 53.40 97.80 99.77 83.80 95.74 Blend 99.34 84.00 100.00 99.55 78.60 97.87 91.23 45.80 93.26 99.08 75.00 98.80 SIG 99.78 89.80 100.00 99.77 87.00 100.00 99.56 35.20 100.00 99.31 68.60 100.00 Adv 100.00 97.20 35.16 100.00 93.40 27.59 100.00 89.20 37.36 99.77 93.80 37.63 Smooth 97.59 86.40 90.11 97.97 85.80 56.38 91.45 28.60 73.63 94.51 62.80 78.02 Nash 98.90 89.80 93.41 98.42 90.00 79.12 1.54 7.00 74.73 94.05 65.20 84.62 Dynamic 100.00 95.00 100.00 100.00 93.40 98.85 100.00 84.60 100.00 100.00 91.00 100.00 WaNet 98.68 84.20 92.05 99.10 84.60 62.64 52.19 14.40 69.32 93.36 58.60 87.50 Average 98.82 90.32 84.72 98.87 88.32 75.97 82.19 50.80 77.20 97.05 76.88 82.70
Appendix B More Experimental Results
B.1 Parallel MTBAs on the ImageNet Subset
Table 6 displays the performance of parallel MTBAs on the ImageNet-20 subset. The experimental results are consistent with that on CIFAR-10. Particularly, most of the triggers can coexist with each other at a high poisoning rate (10%), together achieving a high ASR above 80%. Notably, with an increase in image scale and resolution, the ASR of full-image triggers (Blend, SIG, and Nash) surpasses that of the local triggers (BadNets and Trojan) at a low poisoning rate (1%). We suspect that the full image triggers are somewhat enhanced on high-resolution images, rendering the backdoor features more prominent and easier to learn.
B.2 Hybrid-trigger MTBAs with Different Trigger Stacking Orders
To further explore the cross-activating effect of hybrid-trigger attacks, here we investigated the impact of three different trigger stacking orders with 4 trigger patterns (BadNets, Dynamic, Nash, and WaNet). As shown in Figure 9(a)-(c), we find that 1) the stacking order of the triggers can impact the cross-activating effect. Specifically, the cross-activating effect of a previous trigger may be overwritten or compromised by a subsequent trigger; and 2) the Dynamic trigger exhibits a generally stronger cross-activating capability than others, maintaining high ASR under different stacking orders. Figure 6 provides a few examples of the hybrid triggers attached to the CIFAR-10 images.
B.3 Defense Results on ViTs
Since most existing defense methods were originally designed for CNNs, here we only test the fine-tuning (FT) defense on ViT models. Table 7 reports the defense results of FT against parallel MTBAs (at a poisoning rate of 10%) on the ViT-base model. It is clear that FT is ineffective in defending against parallel MTBAs on ViT, even with different learning rates. The remaining ASRs are still above 80% for 8/12 of the cases. Therefore, how to design effective defense methods for ViT models against parallel MTBAs has become a challenging but imperative task.
B.4 Parallel MTBAs on Different Models
Tables 8, 9, 10, and 11 report the attack performance of parallel MTBAs on the 4 model architectures (2 CNNs and 2 ViTs) at a poisoning rate of 10%, 1%, 0.5% and 0.2%, respectively.
It shows that most triggers can coexist in different model architectures at a high poisoning rate of 10%, but exhibit varied ASRs when the poisoning rate becomes low, e.g., 0.2%.
Table 8: The ASR (%) of parallel MTBAs at the poisoning rate of 10%. The experiment results were averaged on CIFAR-10. The best results are boldfaced.
Parallel MTBAs (poisoning rate 10%) Label Modification Model Clean Acc. OnePixel BadNets Trojan Blend SIG Adv Smooth Nash Dynamic WaNet ResNet-18 92.88 94.53 100.00 99.34 100.00 99.56 99.00 93.22 96.28 100.00 97.59 MobileNet-V2 93.08 95.13 100.00 99.78 99.12 99.12 100.00 93.58 98.45 100.00 99.55 VIT-Small 98.32 94.75 100.00 99.78 99.78 99.78 100.00 98.25 98.69 100.00 97.12 VIT-Base 98.46 91.76 100.00 100.00 99.55 99.78 100.00 98.66 98.89 100.00 97.77 All2One Average 95.69 94.04 100.00 99.73 99.61 99.56 99.75 95.93 98.08 100.00 98.01 ResNet-18 90.66 76.40 86.00 85.40 70.80 76.60 85.00 71.00 75.80 82.20 75.60 MobileNet-V2 92.96 86.20 92.80 91.40 76.00 83.60 91.80 82.20 84.20 86.80 84.40 VIT-Small 98.26 86.80 98.20 97.60 90.40 95.20 97.20 90.80 94.00 96.80 90.60 VIT-Base 98.54 83.60 98.00 98.40 94.00 96.00 97.10 91.20 93.00 96.80 88.00 All2All Average 95.11 83.25 93.75 93.20 82.80 87.85 92.75 83.80 86.75 90.65 84.65 ResNet-18 92.00 89.01 43.01 100.00 100.00 98.90 47.73 88.17 95.56 100.00 96.51 MobileNet-V2 92.60 89.01 53.76 98.90 100.00 98.90 42.05 88.17 95.56 100.00 98.84 VIT-Small 98.60 87.91 54.84 100.00 98.90 100.00 39.77 91.40 97.78 100.00 96.51 VIT-Base 98.00 89.01 54.84 100.00 100.00 100.00 37.50 92.47 96.67 98.84 95.35 All2Random Average 95.30 88.74 51.61 99.73 99.73 99.45 41.76 90.05 96.39 99.71 96.80
Table 9: The ASR (%) of parallel MTBAs at the poisoning rate of 1%. The experiment results were averaged on CIFAR-10. The best results are boldfaced.
Parallel MTBAs (poisoning rate 1%) Label Modification Model Clean Acc. OnePixel BadNets Trojan Blend SIG Adv Smooth Nash Dynamic WaNet ResNet-18 92.62 70.18 100.00 77.80 99.78 92.38 99.10 46.19 20.63 98.21 76.01 MobileNet-V2 93.16 80.00 100.00 92.73 98.86 94.55 100.00 38.64 20.91 98.41 71.36 VIT-Small 98.54 77.22 100.00 96.10 85.90 80.91 98.90 56.83 61.82 99.57 72.02 VIT-Base 98.46 73.25 100.00 98.68 97.81 99.56 99.00 87.06 81.14 100.00 72.81 All2One Average 95.70 75.16 100.00 91.33 95.59 91.85 99.25 57.18 46.12 99.05 73.05 ResNet-18 92.78 38.40 80.80 57.80 37.60 36.60 80.45 14.00 13.20 55.60 28.00 MobileNet-V2 92.62 50.60 80.60 65.80 36.60 32.80 81.60 13.40 11.80 57.80 27.60 VIT-Small 98.4 17.60 35.60 44.40 14.00 15.60 33.30 4.80 1.80 47.40 4.20 VIT-Base 98.66 45.60 65.00 69.40 30.40 28.60 62.20 15.80 17.80 68.20 10.20 All2All Average 95.62 38.05 65.50 59.35 29.65 28.40 64.39 12.00 11.15 57.25 17.50 ResNet-18 92.10 67.01 49.44 60.87 98.88 91.95 44.44 32.97 49.44 96.51 41.11 MobileNet-V2 92.70 68.04 88.76 82.61 98.88 81.61 6.67 38.46 29.21 93.02 62.22 VIT-Small 98.40 59.79 59.55 95.65 68.54 100.00 28.89 38.46 62.92 100.00 42.22 VIT-Base 98.90 71.13 13.48 95.65 87.64 82.76 75.56 42.86 71.91 98.84 48.89 All2Random Average 95.52 66.49 52.81 83.70 88.48 89.08 38.89 38.19 53.37 97.09 48.61
Table 10: The ASR (%) of parallel MTBAs at the poisoning rate of 0.5%. The experiment results were averaged on CIFAR-10. The best results are boldfaced.
Parallel MTBAs (poisoning rate 0.5%) Label Modification Model Clean Acc. OnePixel BadNets Trojan Blend SIG Adv Smooth Nash Dynamic WaNet ResNet-18 92.88 52.71 100.00 63.80 99.32 90.05 93.10 8.37 19.00 96.15 42.31 MobileNet-V2 92.14 55.29 100.00 83.15 99.14 76.67 97.20 9.72 8.42 96.98 27.43 VIT-Small 98.54 63.24 98.17 95.89 67.58 76.71 97.11 14.84 13.70 99.09 14.16 VIT-Base 98.56 37.25 87.36 86.92 62.31 76.72 83.63 13.97 29.27 99.56 8.87 All2One Average 95.53 52.12 96.38 82.44 82.09 80.04 92.76 11.72 17.60 97.94 23.19 ResNet-18 92.82 25.00 62.20 26.00 32.20 25.80 60.12 4.60 1.40 32.80 12.00 MobileNet-V2 92.66 22.60 43.60 33.80 23.40 15.40 41.30 4.60 1.40 32.00 7.60 VIT-Small 98.7 0.00 0.20 3.00 3.80 6.60 0.20 1.60 2.10 13.80 0.60 VIT-Base 98.74 0.40 7.00 18.00 8.00 8.20 6.90 2.80 0.40 20.40 1.20 All2All Average 95.73 12.00 28.25 20.20 16.85 14.00 27.13 3.40 0.85 24.75 5.35 ResNet-18 93.20 34.88 42.70 37.23 83.33 89.47 45.98 7.37 3.45 92.39 2.20 MobileNet-V2 91.50 55.81 0.00 67.02 92.86 77.89 100.00 9.47 5.75 94.57 27.47 VIT-Small 98.20 0.00 1.12 47.87 53.57 69.47 0.00 15.79 4.60 78.26 9.89 VIT-Base 99.10 0.00 29.21 76.60 53.57 73.68 19.54 21.05 26.44 89.13 6.59 All2Random Average 95.50 22.67 18.26 57.18 70.83 77.63 41.38 13.42 10.06 88.59 11.54
Table 11: The ASR (%) of parallel MTBAs at the poisoning rate of 0.2%. The experiment results were averaged on CIFAR-10. The best results are boldfaced.
Parallel MTBAs (poisoning rate 0.2%) Labeling Mode Model Clean Acc. OnePixel BadNets Trojan Blend SIG Adv Smooth Nash Dynamic WaNet ResNet-18 93.62 27.80 99.33 21.52 80.94 70.40 93.13 3.36 3.36 88.57 8.97 MobileNet-V2 93.02 2.51 25.74 6.15 90.89 49.20 20.14 1.37 5.24 86.79 10.48 VIT-Small 98.52 0.23 0.46 53.23 17.28 3.69 0.26 1.61 0.23 88.02 0.69 VIT-Base 98.60 12.02 45.80 54.65 20.41 76.64 40.10 0.91 0.91 97.73 1.59 All2One Average 95.94 10.64 42.83 33.89 52.38 49.98 38.40 1.81 2.43 90.28 5.43 ResNet-18 92.38 1.40 2.20 1.40 19.00 15.20 0.20 4.20 1.00 7.80 3.80 MobileNet-V2 92.46 1.40 1.80 1.60 11.20 10.60 1.10 3.00 1.60 11.40 4.00 VIT-Small 98.54 0.20 0.20 0.40 2.80 1.00 1.10 1.00 0.40 0.20 0.80 VIT-Base 98.64 0.00 0.00 0.00 1.80 1.20 0.00 0.00 0.00 1.80 0.20 All2All Average 95.51 0.75 1.05 0.85 8.70 7.00 0.60 2.05 0.75 5.30 2.20 ResNet-18 93.70 18.39 21.51 14.61 75.53 63.22 53.93 4.49 1.11 83.33 3.26 MobileNet-V2 93.60 31.03 100.00 11.24 92.55 55.17 0.00 1.12 1.11 37.78 6.52 VIT-Small 98.40 0.00 0.00 0.00 14.89 22.99 0.00 0.00 0.00 33.33 2.17 VIT-Base 98.90 0.00 7.53 23.60 10.64 77.01 0.00 0.00 0.00 61.11 2.17 All2Random Average 96.15 12.36 32.26 12.36 48.40 54.60 13.48 1.40 0.56 53.89 3.53

### v4

@@ V-D Re-evaluating Existing Defenses
Backdoor Model Detection. Fig. 7(c) presents the detection AUROC of 4 advanced backdoor model detection methods, including NC, MMBD, UMD, and RNP-U. As can be observed, the four detection methods perform well on All2One multi-trigger attacks, yet perform poorly on All2All and All2Random attacks. Specifically, the detection AUROC deteriorates by nearly 50% and 40% on All2All and All2Random attacks respectively, when compared to that on All2One attacks. Such a huge degradation challenges the reliability of the four detection models under more complex scenarios with unfixed adversarial targets. Out of the four methods, RNP-U consistently outperforms the other three methods, which validates the usefulness of unlearning the benign features before detecting the backdoor features. Meanwhile, although these detection methods can detect the backdoor models to some extent, they all fail to identify the backdoor labels of All2All and All2Random attacks.
Why do MTBAs bypass the existing backdoor detection methods? The primary issue with existing backdoor detection methods, such as NC, is their inability to effectively infer backdoor labels in our all-to-all (all2all) and all-to-random (all2rand) settings. In single-trigger scenarios, a significant statistical difference (e.g., absolute mean deviation) exists between backdoor labels and normal labels, allowing methods like NC to identify backdoor labels based on this statistical deviation. However, in MTBA settings, accurately inferring the true backdoor label becomes challenging when all labels are backdoored.
To validate this, we calculated the minimal perturbation value, i.e., L1-norm, required for most NC-like detection methods under both all2all and all2rand settings in parallel MTBAs. The backdoored ResNet-18 model was trained on CIFAR-10 with a 10% poisoning rate. As illustrated in Fig. 8, the perturbation values across all classes exhibit inconsistency and randomness, making it difficult for NC to accurately infer the true backdoor label category.
TABLE VI: The ASR/CA (%) comparison between single-trigger attacks and parallel MTBAs. The results were obtained on CIFAR-10 with a poisoning rate of 10%.
Backdoor Attacks Metrics OnePixel BadNets Trojan Blend SIG Adv Smooth Nash Dynamic WaNet Average Single-trigger CA 92.69 93.7 92.99 93.7 93 93.19 91.69 93.66 92.99 93.51 93.11 ASR 96.62 100 98.18 100 92.34 99.75 95.56 98.12 99.97 99.1 97.96 Parallel MTBAs CA 92.88 92.88 92.88 92.88 92.88 92.88 92.88 92.88 92.88 92.88 92.88 ASR 94.53 100 99.34 100 99.56 99 93.22 96.28 100 97.59 97.95
@@
Backdoor Removal. Table V reports the defense performance of four backdoor removal methods, including FT, FP, NAD, ANP and ABM. We measure the defense performance by the remaining ASR, the lower the better. None of the defense methods can fully remove the multi-trigger from the model. The average ASRs are still well above 50%, except for NAD and ANP against All2One attacks. Compared to NAD and ANP, fine-tuning and fine-pruning methods FT and FP are essentially ineffective against multi-trigger attacks, with an ASR $60\%$. This is because the presence of multiple triggers prevents effective forgetting of the backdoor correlation during standard fine-tuning. Additionally, the coexistence of multiple triggers leads to inconsistent parameter activations, rendering the activation-based pruning method FP ineffective. The current state-of-the-art defense method ANP performs poorly on All2All and All2Random attacks, limiting its practicability against more realistic attacks that have different target labels. Clean accuracy (CA) is another perspective looking into the effectiveness of the defense method.
The four backdoor removal methods cause a significant reduction in the model’s clean accuracy when facing multi-trigger attacks. Specifically, against All2One attacks, FP, NAD, and ANP reduce the model’s CA by 10.98%, 8.48%, and 8.68% respectively, significantly impairing the functionality of the model. ABM reduces ASR to about 58.7% on All2All while preserving clean accuracy above 89%. However, its effectiveness is still far from sufficient under complex multi-trigger settings. Based on these results, we conclude that existing backdoor detection and removal methods struggle to address the threat of multi-trigger attacks, and developing more advanced countermeasures is imperative.
TABLE VII: The ASR (%) of parallel MTBAs across 4 different model architectures, i.e. ResNet-18, MobileNet-V2, VIT-Small, and VIT-Base, with the total/trigger-wise poisoning rate of 10% (1%). The experiment results were averaged on CIFAR-10. The best results are boldfaced.
Parallel MTBAs (Poisoning rate 10% (1%)) Label Modification Model Clean Acc. OnePixel BadNets Trojan Blend SIG Adv Smooth Nash Dynamic WaNet ResNet-18 92.88 94.53 100.00 99.34 100.00 99.56 99.00 93.22 96.28 100.00 97.59 MobileNet-V2 93.08 95.13 100.00 99.78 99.12 99.12 100.00 93.58 98.45 100.00 99.55 VIT-Small 98.32 94.75 100.00 99.78 99.78 99.78 100.00 98.25 98.69 100.00 97.12 VIT-Base 98.46 91.76 100.00 100.00 99.55 99.78 100.00 98.66 98.89 100.00 97.77 All2One Average 95.69 94.04 100.00 99.73 99.61 99.56 99.75 95.93 98.08 100.00 98.01 ResNet-18 90.66 76.40 86.00 85.40 70.80 76.60 85.00 71.00 75.80 82.20 75.60 MobileNet-V2 92.96 86.20 92.80 91.40 76.00 83.60 91.80 82.20 84.20 86.80 84.40 VIT-Small 98.26 86.80 98.20 97.60 90.40 95.20 97.20 90.80 94.00 96.80 90.60 VIT-Base 98.54 83.60 98.00 98.40 94.00 96.00 97.10 91.20 93.00 96.80 88.00 All2All Average 95.11 83.25 93.75 93.20 82.80 87.85 92.75 83.80 86.75 90.65 84.65 ResNet-18 92.00 89.01 43.01 100.00 100.00 98.90 47.73 88.17 95.56 100.00 96.51 MobileNet-V2 92.60 89.01 53.76 98.90 100.00 98.90 42.05 88.17 95.56 100.00 98.84 VIT-Small 98.60 87.91 54.84 100.00 98.90 100.00 39.77 91.40 97.78 100.00 96.51 VIT-Base 98.00 89.01 54.84 100.00 100.00 100.00 37.50 92.47 96.67 98.84 95.35 All2Random Average 95.30 88.74 51.61 99.73 99.73 99.45 41.76 90.05 96.39 99.71 96.80
V-E Analysis and Further Exploration
Single-Trigger vs. Multi-Trigger Attacks. Table VI presents the comparative performance of single-trigger and multi-trigger (Parallel MTBA) attacks on a CIFAR-10 dataset with a ResNet-18 model and a consistent poisoning rate of 10%. The results indicate that multi-trigger attacks achieve nearly identical Attack Success Rates (ASR) and Clean Accuracy (CA) compared to single-trigger attacks. For example, the average CA for single-trigger attacks is 93.11%, while it is 92.88% for multi-trigger attacks, showing only a 0.23% difference. Similarly, the average ASR for single-trigger attacks is 97.96%, compared to 97.95% for multi-trigger attacks, reflecting a negligible 0.01% difference. These values suggest that introducing multiple triggers does not significantly degrade the model’s performance on clean data or the effectiveness of the attack.
In terms of specific cases, single-trigger attacks with the BadNets method achieve an ASR of 100% and a CA of 93.7%, while multi-trigger attacks with BadNets also reach an ASR of 100% with a slightly lower CA of 92.88%. Another example is the Dynamic attack, where the single-trigger ASR is 99.97% and the CA is 92.99%, closely matched by the multi-trigger ASR of 100% and CA of 92.88%. These similarities across various methods reinforce that multi-trigger attacks can maintain high ASR and CA values, effectively balancing attack strength and model utility.
The minimal impact on CA suggests that models compromised with multiple triggers remain highly functional for benign tasks, making detection more challenging. The ability of multi-trigger attacks to maintain similar ASR to single-trigger attacks, despite involving multiple triggers, highlights their robustness and adaptability across diverse attack scenarios. This increased attack diversity allows multi-trigger methods to be more flexible and resilient, potentially posing a more significant threat in real-world applications where detection is crucial.
Parallel MTBAs Across different Architectures. Table VII reports the attack performance of parallel MTBAs on four distinct model architectures, including two Convolutional Neural Networks (CNNs) (i.e. ResNet-18 and MobileNet-V2) and two Vision Transformers (ViTs) (i.e. VIT-Small and VIT-Base), at a poisoning rate of 10%. The results indicate that MTBAs achieve high ASR and maintain CA across different architectures, demonstrating a high degree of generalizability.
Under the All2One label modification, ResNet-18 achieves a clean accuracy of 92.88% with ASR values of 100% for both BadNets and Dynamic attacks. Similarly, VIT-Base achieves a clean accuracy of 98.46% and reaches 100% ASR with the BadNets, Trojan, and Dynamic methods. These high ASR values across both CNN and ViT architectures suggest that parallel MTBAs are highly effective across different model structures without significantly impacting clean accuracy, with an average CA of 95.69% for All2One label modification. However, the effectiveness of certain attacks varies slightly based on architecture. For example, in the All2All label modification, ResNet-18 shows lower ASR values for attacks like OnePixel (76.40%) and SIG (76.60%), compared to MobileNet-V2, which achieves higher ASR values of 86.20% for OnePixel and 83.60% for SIG. This variation may be attributed to differences in architectural sensitivity to small perturbations, as CNNs and ViTs process input data differently, with ViTs relying on token-based attention mechanisms that may slightly reduce the influence of certain types of injected triggers.
These findings highlight the versatility of MTBAs in compromising various architectures while maintaining high clean accuracy, especially at higher poisoning rates. The consistent effectiveness of attacks across architectures underscores the need for adaptive defense mechanisms that can detect and mitigate MTBAs in both CNNs and ViTs, accommodating differences in architectural vulnerabilities and the persistence of high ASR values.
Hybrid-Trigger MTBAs with Different Stacking Orders. To further explore the impact of hybrid-trigger attacks, we analyzed how different stacking orders of four trigger patterns—BadNets, Dynamic, Nash, and WaNet—affect their cross-activation capabilities. As shown in Figures 9(a)-(c), our analysis reveals two important observations: 1) the order of trigger stacking influences the overall cross-activating effect, where triggers applied earlier may be compromised or overwritten by subsequent ones, diminishing their effectiveness; and 2) the Dynamic trigger consistently demonstrates stronger cross-activating capabilities compared to the other patterns, maintaining a high ASR across various stacking sequences. This indicates that certain triggers, such as Dynamic, possess intrinsic properties that make them more resistant to interference from other triggers, thereby sustaining their attack potential. These findings emphasize the importance of trigger selection and ordering in hybrid-trigger attacks, as the interplay between triggers can significantly impact the overall effectiveness of the attack.
These hybrid triggers present a significant challenge for traditional backdoor detection methods, as the combination of multiple trigger patterns complicates both their identification and mitigation. This underscores the need for advanced defense mechanisms that can handle complex, multi-trigger backdoor attacks and adapt to their varying stacking strategies. In summary, the analysis highlights the sophistication and adaptability of multi-trigger backdoor attacks, as well as the critical need for robust, architecture-aware defenses to counteract them effectively.
VI Discussion
This work introduces three poisoning strategies for multi-trigger backdoor attacks (MTBAs): parallel, sequential, and hybrid, covering scenarios involving both independent and collusive adversaries. Beyond the empirical findings, several important insights emerge that can guide future research on understanding and defending against MTBAs.
VI-1 Potential Directions for Hybrid Defense Mechanisms
Existing defenses are typically designed to operate at a single stage of the backdoor pipeline, such as backdoor unlearning or backdoor removal. However, combining techniques across stages may offer stronger protection. We conducted preliminary evaluations that integrate backdoor unlearning (BU) and removal-based methods (BR). Results show that hybrid approaches—such as combining BU with fine-tuning (BU + FT), adversarial neuron pruning (BU + ANP), or anomaly-based learning (BU + ABL)—achieve significantly lower attack success rates (ASR $\leq$ 3%) while preserving clean accuracy above 90%. This suggests that chaining complementary defenses, e.g., exposure via unlearning followed by backdoor removal, can be more effective than using any single method alone.
VI-2 Towards a Theoretical Understanding of Trigger Interactions
While our work primarily emphasizes empirical evaluation, a deeper theoretical understanding of multi-trigger attacks remains both critical and underexplored in backdoor research. We highlight three preliminary perspectives that shed light on why and how triggers interact: (1) Representational dominance, where salient triggers with stronger and more visible features (e.g., BadNets, Dynamic) dominate the optimization process, thereby overwriting weaker and stealthier triggers (e.g., Nash, WaNet); (2) Shared feature representation, where multiple triggers overlap in representational subspaces, causing their gradient signals to interfere and jointly activate shared neurons, leading to cross-activation effects; and (3) Parametric memory, where residual traces of previously learned triggers persist in the weight space even after being suppressed, allowing them to partially re-emerge when reinforced by correlated patterns.
Providing a rigorous theoretical analysis of these mechanisms is both important and challenging, and we acknowledge that this aspect remains underexplored in the backdoor literature. As part of our future work, we plan to investigate these dynamics in greater depth using information-theoretic and spectral tools (e.g., mutual information, singular value analysis), with the goal of establishing a more principled foundation for understanding multi-trigger backdoors.
VII Conclusion